# Infinity categories with duality and hermitian multiplicative infinite loop space machines

### v3

@@ 3 Hermitian objects
which is an equivalence by definition of the join construction. ∎
@@
Let $(\mathrm{X},\tau)$ be space with $C_{2}$-action. There is a canonical equivalence over X:normal-:normal-$\mathrm{X}:$
@@ 4 Direct sum real $K$-theory
  • Evaluation at *$*\in\mathsf{Fin}[C_{2}]\to\mathrm{Span}(\mathsf{Fin}[C_{2}])$ induces a functor $(-)^{C_{2}}:\mathrm{Sp}^{C_{2}}\to\mathrm{Sp}$.
@@
By [21, Theorem 2.2.6.2., Proposition 2.2.6.16.] for every small $\infty$-category ${\mathcal{C}}$ and presentably symmetric monoidal $\infty$-category ${\mathcal{D}}$ there is a presentably symmetric monoidal structure on the $\infty$-category $\mathrm{Fun}({\mathcal{C}},{\mathcal{D}})$ given by Day-convolution characterized by the following universal property: the evaluation functor $\mathrm{Fun}({\mathcal{C}},{\mathcal{D}})\times{\mathcal{C}}\to{\mathcal{D}}$ is symmetric monoidal and for every symmetric monoidal $\infty$-category ${\mathcal{B}}$ the canonical functor
@@ 5 Symmetric monoidal $\infty$-categories with duality
Since the opposite of a rigid symmetric monoidal $\infty$-category is rigid, the non-trivial $C_{2}$-action on $\mathrm{Cmon}(\mathsf{Cat}_{\infty})$ restricts to $\mathrm{Cmon}(\mathsf{Cat}_{\infty})^{\mathrm{rig}}$ and $\mathrm{Cmon}(\mathsf{Cat}_{\infty})^{\mathrm{rig}}_{\mathrm{preadd}}.$
@@
A functor ${\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}^{\mathrm{op}}\to{\mathcal{S}}$ is classified via the straightening equivalence [22, Theorem 3.2.0.1.] by a right fibration ${\mathcal{D}}\to{\mathcal{C}}\times{\mathcal{C}}$. A $C_{2}$-equivariant functor $\widetilde{{\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}^{\mathrm{op}}}\to{\mathcal{S}}$ is classified by a $C_{2}$-equivariant right fibration ${\mathcal{D}}\to\widetilde{{\mathcal{C}}\times{\mathcal{C}}}$. A $C_{2}$-equivariant functor $\widetilde{{\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}^{\mathrm{op}}}\to{\mathcal{S}}$ corresponds to a functor $(\widetilde{{\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}^{\mathrm{op}}})_{hC_{2}}\to{\mathcal{S}}$ that is classified by a right fibration ${\mathcal{D}}\to(\widetilde{{\mathcal{C}}\times{\mathcal{C}}})_{hC_{2}}$.
@@
The $C_{2}$-action on $\mathsf{Cat}_{\infty}$ gives rise to a $C_{2}$-action on $\mathrm{Fun}([1],\mathsf{Cat}_{\infty})$ such that the functor $\rho:\mathrm{Fun}([1],\mathsf{Cat}_{\infty})\to\mathsf{Cat}_{\infty}$ evaluating at the target is $C_{2}$-equivariant. Thus the pullback $\mathsf{Cat}_{\infty/{\mathcal{C}}}$ of the $C_{2}$-equivariant functor $\rho$ along the $C_{2}$-equivariant functor *$*\to\mathsf{Cat}_{\infty}$ corresponding to the homotopy $C_{2}$-fixed point ${\mathcal{C}}$ of $\mathsf{Cat}_{\infty}$ carries a $C_{2}$-action and $(\mathsf{Cat}_{\infty/{\mathcal{C}}})^{hC_{2}}$ is the fiber of the induced functor $\rho^{hC_{2}}:\mathrm{Fun}([1],\mathsf{Cat}_{\infty})^{hC_{2}}\to(\mathsf{Cat}_{\infty},(-)^{\mathrm{op}})^{hC_{2}}$ over ${\mathcal{C}}.$ The functor $\rho^{hC_{2}}$ identifies with the functor $\mathrm{Fun}([1],(\mathsf{Cat}_{\infty},(-)^{\mathrm{op}})^{hC_{2}})\to(\mathsf{Cat}_{\infty},(-)^{\mathrm{op}})^{hC_{2}}$ evaluating at the target whose fiber over ${\mathcal{C}}$ is $((\mathsf{Cat}_{\infty},(-)^{\mathrm{op}})^{hC_{2}})_{/{\mathcal{C}}}.$
@@
We prove the following theorem, for which we use the twisted arrow right fibration $\mathrm{Tw}({\mathcal{C}})\to{\mathcal{C}}\times{\mathcal{C}}^{\mathrm{op}}$ associated to any $\infty$-category ${\mathcal{C}},$ which by straightening classifies the mapping space functor ${\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}\to{\mathcal{S}}$ [22, Proposition 5.2.1.3.].
@@
Since $BC_{2}$ is a space, straightening [22, Theorem 3.2.0.1.] gives an equivalence $\mathsf{Cat}_{\infty}[C_{2}]\simeq\mathsf{Cat}_{\infty/BC_{2}}.$ Hence we have a right fibration $\mathsf{Cat}_{\infty}[C_{2}]\simeq\mathsf{Cat}_{\infty/BC_{2}}\to\mathsf{Cat}_{\infty}$ that sends a small $\infty$-category with $C_{2}$-action ${\mathcal{D}}$ classified by a functor $X\to BC_{2}$ to $X$, where $X$ has the universal property of the coinvariants ${\mathcal{D}}_{hC_{2}}$, the colimit of the functor $BC_{2}\to\mathsf{Cat}_{\infty}$ encoding the $C_{2}$-action on ${\mathcal{D}}$, by [22, Corollary 3.3.4.3.]. The right fibration $\mathsf{Cat}_{\infty}[C_{2}]\to\mathsf{Cat}_{\infty}$ yields an equivalence
@@
over $\widetilde{\mathsf{Cat}_{\infty}\times\mathsf{Cat}_{\infty}}.$ By [13, Theorem 1] the functor $\zeta$ is an equivalence.
@@
where the latter is the Yoneda-embedding. The functor $(-)\otimes A:{\mathcal{C}}\to{\mathcal{C}}$ is an equivalence because $A$ is tensor-invertible. The functor $(-)^{\vee}:{\mathcal{C}}^{\mathrm{op}}\to{\mathcal{C}}$ is an equivalence by [14, Lemma 3.11.].
@@
The latter corresponds to a map $\alpha^{\mathrm{op}}_{!}({\mathcal{C}}(-,\mathrm{X}^{\vee}\otimes A))\to{\mathcal{D}}(-,\alpha(X)^{\vee}\otimes\alpha(A))$, which is an equivalence by [22, Proposition 5.2.6.3.].
@@
[…]
@@ 6 Linear dualities
The $\infty$-category ${\mathrm{Mod}}_{{\mathcal{C}}}(\mathsf{Cat}_{\infty})$ carries a canonical closed action of $\mathsf{Cat}_{\infty}$ and so is enriched in $\mathsf{Cat}_{\infty}$ [15, Proposition 6.10. ], what one calls an $(\infty,2)$-category. Hence the full subcategory $B{\mathcal{C}}$ may be viewed as an $(\infty,2)$-category as well.
@@
Every symmetric monoidal functor ${\mathcal{C}}\to{\mathcal{D}}$ gives rise to a symmetric monoidal functor ${\mathcal{D}}\otimes_{{\mathcal{C}}}(-):{\mathrm{Mod}}_{{\mathcal{C}}}(\mathsf{Cat}_{\infty})\to{\mathrm{Mod}}_{{\mathcal{D}}}(\mathsf{Cat}_{\infty})$ that restricts to a symmetric monoidal functor $B{\mathcal{C}}\to B{\mathcal{D}}.$
@@
In the following we use that a symmetric monoidal $\infty$-category with $\mathrm{G}$-action ${\mathcal{B}}$, i.e. a functor $B\mathrm{G}\to\mathrm{Cmon}(\mathsf{Cat}_{\infty})$ is classified by a $B\mathrm{G}$-family of $\infty$-operads ${\mathcal{B}}^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$ in the sense of [21, Definition 2.3.2.10.], where the fiber over the object of $B\mathrm{G}$ is a symmetric monoidal $\infty$-category. By functoriality of $\mathrm{Calg},{\mathrm{Mod}}$ the $B\mathrm{G}$-family of $\infty$-operads ${\mathcal{D}}^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$ whose fiber over the object of $B\mathrm{G}$ is a symmetric monoidal $\infty$-category, induces $B\mathrm{G}$-families of $\infty$-operads $\mathrm{Calg}({\mathcal{D}})^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$, ${\mathrm{Mod}}({\mathcal{D}})^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$ whose fiber over the object of $B\mathrm{G}$ is a symmetric monoidal $\infty$-category.
@@
1.: The symmetric monoidal $\infty$-category $\mathrm{Calg}({\mathcal{D}}^{h\mathrm{G}})^{\otimes}\to\mathsf{Fin}_{*}$ is cocartesian [21, Proposition 3.2.4.7.]. So by [21, Proposition 2.4.3.9.] the object $A\in\mathrm{Calg}({\mathcal{D}}^{h\mathrm{G}})$ uniquely lifts to a commutative algebra in $\mathrm{Calg}({\mathcal{D}}^{h\mathrm{G}})$, i.e. a lax symmetric monoidal functor $\mathsf{Fin}_{*}\to\mathrm{Calg}({\mathcal{D}}^{h\mathrm{G}})^{\otimes}$. The latter corresponds by adjointness to a map of $B\mathrm{G}$-families of $\infty$-operads $\mathsf{Fin}_{*}\times B\mathrm{G}\to\mathrm{Calg}({\mathcal{D}})^{\otimes}.$ We define a $B\mathrm{G}$-family of $\infty$-operads by the pullback of $B\mathrm{G}$-families of $\infty$-operads:
@@
The right vertical functor is a cocartesian fibration by [21, Theorem 4.5.3.1.], where we use that ${\mathcal{D}}$ admits geometric realizations that are preserved by the tensor product. This implies that ${\mathrm{Mod}}_{A}({\mathcal{D}})^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$ is a cocartesian fibration so that its fiber over the object of $B\mathrm{G}$ is a symmetric monoidal $\infty$-category. Applying the right adjoint $((-)^{h\mathrm{G}})^{\otimes}$ we obtain a pullback square of $\infty$-operads
@@
Every additive $\infty$-category ${\mathcal{D}}$ is canonically enriched in the $\infty$-category $\mathrm{Sp}_{\geq 0}$ of connective spectra. In particular, for every $\ell\geq 0$ and $\mathrm{X}\in{\mathcal{D}}$ the tensor $S^{\oplus\ell}\otimes\mathrm{X}$ of the $\ell$-fold sum of the shere spectrum $S$ and $\mathrm{X}$ is defined, and is given by Xsuperscriptnormal-$\mathrm{X}^{\oplus\ell}$. Moreover there is a canonical equivalence $(S,(-)^{\vee})\otimes(\mathrm{X},\mathrm{triv})\simeq(\mathrm{X},(-)^{\vee})$ in ${\mathcal{D}}[C_{2}]$ lying over the identity and dually, an equivalence $(S,(-)_{\vee})\otimes(\mathrm{X},\mathrm{triv})\simeq(\mathrm{X},(-)_{\vee})$ in ${\mathcal{D}}[C_{2}]$ lying over the identity.

### v4

@@ 3 Hermitian objects
which is an equivalence by definition of the join construction.
@@
Let $(\mathrm{X},\tau)$ be space with $C_{2}$-action. There is a canonical equivalence over $\mathrm{X}:$
@@ 4 Direct sum real $K$-theory
  • Evaluation at $*\in\mathsf{Fin}[C_{2}]\to\mathrm{Span}(\mathsf{Fin}[C_{2}])$ induces a functor $(-)^{C_{2}}:\mathrm{Sp}^{C_{2}}\to\mathrm{Sp}$.
@@
By [25, Theorem 2.2.6.2., Proposition 2.2.6.16.] for every small $\infty$-category ${\mathcal{C}}$ and presentably symmetric monoidal $\infty$-category ${\mathcal{D}}$ there is a presentably symmetric monoidal structure on the $\infty$-category $\mathrm{Fun}({\mathcal{C}},{\mathcal{D}})$ given by Day-convolution characterized by the following universal property: the evaluation functor $\mathrm{Fun}({\mathcal{C}},{\mathcal{D}})\times{\mathcal{C}}\to{\mathcal{D}}$ is symmetric monoidal and for every symmetric monoidal $\infty$-category ${\mathcal{B}}$ the canonical functor
@@ 5 Symmetric monoidal $\infty$-categories with duality
Since the opposite of a rigid symmetric monoidal $\infty$-category is rigid, the non-trivial $C_{2}$-action on $\mathrm{Cmon}(\mathsf{Cat}_{\infty})$ restricts to $\mathrm{Cmon}(\mathsf{Cat}_{\infty})^{\mathrm{rig}}$ and Cmon($\mathrm{Cmon}(\mathsf{Cat}_{\infty})^{\mathrm{rig}}_{\mathrm{preadd}}.$
@@
A functor ${\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}^{\mathrm{op}}\to{\mathcal{S}}$ is classified via the straightening equivalence [26, Theorem 3.2.0.1.] by a right fibration ${\mathcal{D}}\to{\mathcal{C}}\times{\mathcal{C}}$. A $C_{2}$-equivariant functor $\widetilde{{\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}^{\mathrm{op}}}\to{\mathcal{S}}$ is classified by a $C_{2}$-equivariant right fibration ${\mathcal{D}}\to\widetilde{{\mathcal{C}}\times{\mathcal{C}}}$. A $C_{2}$-equivariant functor $\widetilde{{\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}^{\mathrm{op}}}\to{\mathcal{S}}$ corresponds to a functor $(\widetilde{{\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}^{\mathrm{op}}})_{hC_{2}}\to{\mathcal{S}}$ that is classified by a right fibration ${\mathcal{D}}\to(\widetilde{{\mathcal{C}}\times{\mathcal{C}}})_{hC_{2}}$.
@@
The $C_{2}$-action on $\mathsf{Cat}_{\infty}$ gives rise to a $C_{2}$-action on $\mathrm{Fun}([1],\mathsf{Cat}_{\infty})$ such that the functor $\rho:\mathrm{Fun}([1],\mathsf{Cat}_{\infty})\to\mathsf{Cat}_{\infty}$ evaluating at the target is $C_{2}$-equivariant. Thus the pullback $\mathsf{Cat}_{\infty/{\mathcal{C}}}$ of the $C_{2}$-equivariant functor $\rho$ along the $C_{2}$-equivariant functor $*\to\mathsf{Cat}_{\infty}$ corresponding to the homotopy $C_{2}$-fixed point ${\mathcal{C}}$ of $\mathsf{Cat}_{\infty}$ carries a $C_{2}$-action and $(\mathsf{Cat}_{\infty/{\mathcal{C}}})^{hC_{2}}$ is the fiber of the induced functor $\rho^{hC_{2}}:\mathrm{Fun}([1],\mathsf{Cat}_{\infty})^{hC_{2}}\to(\mathsf{Cat}_{\infty},(-)^{\mathrm{op}})^{hC_{2}}$ over ${\mathcal{C}}.$ The functor $\rho^{hC_{2}}$ identifies with the functor $\mathrm{Fun}([1],(\mathsf{Cat}_{\infty},(-)^{\mathrm{op}})^{hC_{2}})\to(\mathsf{Cat}_{\infty},(-)^{\mathrm{op}})^{hC_{2}}$ evaluating at the target whose fiber over ${\mathcal{C}}$ is $((\mathsf{Cat}_{\infty},(-)^{\mathrm{op}})^{hC_{2}})_{/{\mathcal{C}}}.$
@@
We prove the following theorem, for which we use the twisted arrow right fibration $\mathrm{Tw}({\mathcal{C}})\to{\mathcal{C}}\times{\mathcal{C}}^{\mathrm{op}}$ associated to any $\infty$-category ${\mathcal{C}},$ which by straightening classifies the mapping space functor ${\mathcal{C}}^{\mathrm{op}}\times{\mathcal{C}}\to{\mathcal{S}}$ [26, Proposition 5.2.1.3.].
@@
Since $BC_{2}$ is a space, straightening [26, Theorem 3.2.0.1.] gives an equivalence $\mathsf{Cat}_{\infty}[C_{2}]\simeq\mathsf{Cat}_{\infty/BC_{2}}.$ Hence we have a right fibration $\mathsf{Cat}_{\infty}[C_{2}]\simeq\mathsf{Cat}_{\infty/BC_{2}}\to\mathsf{Cat}_{\infty}$ that sends a small $\infty$-category with $C_{2}$-action ${\mathcal{D}}$ classified by a functor $X\to BC_{2}$ to $X$, where $X$ has the universal property of the coinvariants ${\mathcal{D}}_{hC_{2}}$, the colimit of the functor $BC_{2}\to\mathsf{Cat}_{\infty}$ encoding the $C_{2}$-action on ${\mathcal{D}}$, by [26, Corollary 3.3.4.3.]. The right fibration $\mathsf{Cat}_{\infty}[C_{2}]\to\mathsf{Cat}_{\infty}$ yields an equivalence
@@
over $\widetilde{\mathsf{Cat}_{\infty}\times\mathsf{Cat}_{\infty}}.$ By [16, Theorem 5.2.] the functor $\zeta$ is an equivalence.
@@
where the latter is the Yoneda-embedding. The functor $(-)\otimes A:{\mathcal{C}}\to{\mathcal{C}}$ is an equivalence because $A$ is tensor-invertible. The functor $(-)^{\vee}:{\mathcal{C}}^{\mathrm{op}}\to{\mathcal{C}}$ is an equivalence by [13, Lemma 3.10.].
@@
The latter corresponds to a map $\alpha^{\mathrm{op}}_{!}({\mathcal{C}}(-,\mathrm{X}^{\vee}\otimes A))\to{\mathcal{D}}(-,\alpha(X)^{\vee}\otimes\alpha(A))$, which is an equivalence by [26, Proposition 5.2.6.3.].
@@
that assigns to a preadditive rigid symmetric monoidal $\infty$-category ${\mathcal{C}}$ a real $K$-theory ${\mathbf{E}}_{\infty}$-$C_{2}$-spectrum $\mathrm{KR}({\mathcal{C}},(-)^{\vee}).$
@@ 6 Linear dualities
The $\infty$-category ${\mathrm{Mod}}_{{\mathcal{C}}}(\mathsf{Cat}_{\infty})$ carries a canonical closed action of $\mathsf{Cat}_{\infty}$ and so is enriched in $\mathsf{Cat}_{\infty}$ [14, Proposition 6.10.] or [17, Example 3.23.], what one calls an $(\infty,2)$-category. Hence the full subcategory $B{\mathcal{C}}$ may be viewed as an $(\infty,2)$-category as well.
@@
Every symmetric monoidal functor ${\mathcal{C}}\to{\mathcal{D}}$ gives rise to a symmetric monoidal functor ${\mathcal{D}}\otimes_{{\mathcal{C}}}(-):{\mathrm{Mod}}_{{\mathcal{C}}}(\mathsf{Cat}_{\infty})\to{\mathrm{Mod}}_{{\mathcal{D}}}(\mathsf{Cat}_{\infty})$ that restricts to a symmetric monoidal functor $B{\mathcal{C}}\to B{\mathcal{D}}.$ The latter is the one-object case of $\infty\mathsf{Cat}$-enriched Yoneda-extension (see [21] or [15] for enriched presheaves and [heine2024higher, Corollary 5.27.] for enriched Day-convolution).
@@
In the following we use that a symmetric monoidal $\infty$-category with $\mathrm{G}$-action ${\mathcal{B}}$, i.e. a functor $B\mathrm{G}\to\mathrm{Cmon}(\mathsf{Cat}_{\infty})$ is classified by a $B\mathrm{G}$-family of $\infty$-operads ${\mathcal{B}}^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$ in the sense of [25, Definition 2.3.2.10.], where the fiber over the object of $B\mathrm{G}$ is a symmetric monoidal $\infty$-category. By functoriality of $\mathrm{Calg},{\mathrm{Mod}}$ the $B\mathrm{G}$-family of $\infty$-operads ${\mathcal{D}}^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$ whose fiber over the object of $B\mathrm{G}$ is a symmetric monoidal $\infty$-category, induces $B\mathrm{G}$-families of $\infty$-operads $\mathrm{Calg}({\mathcal{D}})^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$, ${\mathrm{Mod}}({\mathcal{D}})^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$ whose fiber over the object of $B\mathrm{G}$ is a symmetric monoidal $\infty$-category.
@@
1.: The symmetric monoidal $\infty$-category $\mathrm{Calg}({\mathcal{D}}^{h\mathrm{G}})^{\otimes}\to\mathsf{Fin}_{*}$ is cocartesian [25, Proposition 3.2.4.7.]. So by [25, Proposition 2.4.3.9.] the object $A\in\mathrm{Calg}({\mathcal{D}}^{h\mathrm{G}})$ uniquely lifts to a commutative algebra in $\mathrm{Calg}({\mathcal{D}}^{h\mathrm{G}})$, i.e. a lax symmetric monoidal functor $\mathsf{Fin}_{*}\to\mathrm{Calg}({\mathcal{D}}^{h\mathrm{G}})^{\otimes}$. The latter corresponds by adjointness to a map of $B\mathrm{G}$-families of $\infty$-operads $\mathsf{Fin}_{*}\times B\mathrm{G}\to\mathrm{Calg}({\mathcal{D}})^{\otimes}.$ We define a $B\mathrm{G}$-family of $\infty$-operads by the pullback of $B\mathrm{G}$-families of $\infty$-operads:
@@
The right vertical functor is a cocartesian fibration by [25, Theorem 4.5.3.1.], where we use that ${\mathcal{D}}$ admits geometric realizations that are preserved by the tensor product. This implies that ${\mathrm{Mod}}_{A}({\mathcal{D}})^{\otimes}\to\mathsf{Fin}_{*}\times B\mathrm{G}$ is a cocartesian fibration so that its fiber over the object of $B\mathrm{G}$ is a symmetric monoidal $\infty$-category. Applying the right adjoint $((-)^{h\mathrm{G}})^{\otimes}$ we obtain a pullback square of $\infty$-operads
@@
Every additive $\infty$-category ${\mathcal{D}}$ is canonically enriched in the $\infty$-category $\mathrm{Sp}_{\geq 0}$ of connective spectra. In particular, for every $\ell\geq 0$ and $\mathrm{X}\in{\mathcal{D}}$ the tensor $S^{\oplus\ell}\otimes\mathrm{X}$ of the $\ell$-fold sum of the shere spectrum $S$ and $\mathrm{X}$ is defined, and is given by $\mathrm{X}^{\oplus\ell}$. Moreover there is a canonical equivalence $(S,(-)^{\vee})\otimes(\mathrm{X},\mathrm{triv})\simeq(\mathrm{X},(-)^{\vee})$ in ${\mathcal{D}}[C_{2}]$ lying over the identity and dually, an equivalence $(S,(-)_{\vee})\otimes(\mathrm{X},\mathrm{triv})\simeq(\mathrm{X},(-)_{\vee})$ in ${\mathcal{D}}[C_{2}]$ lying over the identity.